# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ E.2 Brickwork circuits
For brickwork circuits of depth $d_{C}=4$, we use the fact that we can compile layers 1 and 2 together, as well as layers 3 and 4, which results in a circuit of the same architecture as the depth-$d_{C}=2$ 4-local circuits (see Fig. 3). Accordingly, our approach is very similar to before: We use cells of size $8\times 4$, and absorb the final two layers of gates (which all act within single cells) into the local measurement channels $\mathcal{T}^{Q_{i}\rightarrow M_{i}}$. The decomposition Eq. (83) still holds, but this time, the maps $\Theta_{i,a}$ will not correspond to local 2-design measurements on each quartet of qudits, since the underlying random unitaries are 2-local rather than 4-local. For this reason, we split up each quartet further into two pairs of qudits $a=(b_{1},b_{2})$, such that the qudits in each pair $b_{1,2}$ are acted on by the same unitary in the final layer of gates. We claim that the map $\tilde{\Theta}_{i,a}$ for each quartet satisfies
@@
where $X_{b_{1}}=\operatorname{Tr}_{b_{2}}[X_{a}]$ is the result of a partial trace on $b_{2}$, and similar for $X_{b_{2}}$. Evidently, for large $q$, these latter contributions which were not present before will be suppressed.
@@
where $\mathcal{P}_{Y}\coloneqq\text{id}_{Y}-\mathcal{D}_{Y}$ is a map that projects onto traceless operators on some region $Y$. Our aim is to compute $\|\tilde{\Theta}_{i,a}[X]\|_{2}$ for an arbitrary operator $X$, so we can employ the triangle inequality to bound this by the sum of three contributions corresponding to the three terms in the above expansion.
Recall that the channel $\mathcal{T}_{i,a}$ includes the action of the random unitaries in the third and fourth layers of gates which act within the quartet $a$, along with the projective measurements. This can be represented as a generalized measurement, and hence we can use the results from Appendix C, specifically Eq. (76). Specifically, using the definition of the map $\Phi$ in Eq. (77), we know that there is a choice of $\sigma_{i,a}$ such that
[…]
and similarly for the other two components of $\Theta_{i,a}$ in Eq. (91), with the maps $\mathcal{P}_{i,b_{1}}$, $\mathcal{P}_{i,b_{2}}$ replaced with $\mathcal{D}_{b_{1}}$, $\mathcal{D}_{b_{2}}$ as appropriate. To compute the map $\Phi$, we need the second moments of the POVM ensemble, which in this instance features states $U^{\dagger}\ket{m}\!\!\bra{m}U$ occurring with probability $p_{U}/d_{b_{1}b_{2}}$, where $\{(p_{U},U)\}$ is the ensemble of pre-measurement unitaries applied to the quartet $a$. The unitaries $U$ are made up of four two-qudit gates, all of which are sampled from a 2-design: The first two gates applied in parallel connect $b_{1}$ and $b_{2}$, while the last two gates act independently on $b_{1}$ and $b_{2}$. Using the second moments of the Haar distribution over $\mathrm{U}(q^{2})$, we can compute
$\Phi=\frac{q^{8}}{(q^{4}-1)^{2}}\sum_{\tau_{b_{1}}\tau_{b_{2}}}\left(\frac{2q}{q^{2}+1}\right)^{2(1-\delta_{\tau_{1},\tau_{2}})}\Upsilon_{\tau_{b_{1}}}\otimes\Upsilon_{\tau_{b_{1}}}$ (93)
where we define the maps
$\Upsilon_{\tau}=\begin{dcases}q^{2}\mathcal{D}-\frac{1}{q^{2}}\text{id}&\tau=+1\\ \text{id}-\mathcal{D}&\tau=-1\end{dcases}$ (94)
This can be substituted directly into Eq. (92) to get
@@
where $\tilde{X}=X-\mathcal{D}[X]$ is the traceless part of $X$. For the other terms in Eq. (91), we have
@@ Appendix F Proof of advantage over shallow classical circuits
Finally we use Lemma 7 to upper bound the mean error $\overline{\epsilon}$ in the case where $I$ and $J$ are traced out, and $H$ is sent through the distillation channel. Again we can look at a sum over separating domain walls with Boltzmann weights are given by $2^{-m|W|/2}$, i.e. the effective inverse temperature is $\beta=\frac{1}{2}m\log 2$. In this case, since $H$, $I$, and $J$ are in the bulk of the system, the relevant domain walls now include loop configurations of the following kinds: 1) a single loop encircling $H$, 2) a single loop encircling $I$ and $J$ together, and 3) two separate loops encircling $I$ and $J$ separately. As we will see, configurations with long loops are strongly suppressed. For a $\sqrt{n}\times\sqrt{n}$ square lattice, we will find it useful to condition on the event that $H$, $I$, and $J$ are all a $\ell_{1}$ distance of at least $\sqrt{n}/4$ from the edges of the system, and also all pairwise separated by at least $\sqrt{n}/4$ (this occurs with constant probability). This effectively allows us to ignore configurations of type 2) above, and also any configurations with domain walls that terminate on the edges of the system (and thus do not form closed loops).

### v2

@@ E.2 Brickwork circuits
For brickwork circuits of depth $d_{C}=4$, we use the fact that we can compile layers 1 and 2 together, as well as layers 3 and 4, which results in a circuit of the same architecture as the depth-$d_{C}=2$ 4-local circuits (see Fig. 3). Accordingly, our approach is very similar to before: We use cells of size $8\times 4$, and absorb the final two layers of gates (which all act within single cells) into the local measurement channels $\mathcal{T}^{Q_{i}\rightarrow M_{i}}$. The decomposition Eq. (85) still holds, but this time, the maps $\Theta_{i,a}$ will not correspond to local 2-design measurements on each quartet of qudits, since the underlying random unitaries are 2-local rather than 4-local; thus Eq. (57) no longer holds. For this reason, we split up each quartet further into two pairs of qudits $a=(b_{1},b_{2})$, such that the qudits in each pair $b_{1,2}$ are acted on by the same unitary in the final layer of gates. We claim that the conjugated map $\tilde{\Theta}_{i,a}$ for each quartet satisfies
@@
where $X_{b_{1}}=\operatorname{Tr}_{b_{2}}[X_{a}]$ is the result of a partial trace on $b_{2}$, and similar for $X_{b_{2}}$. Evidently, for large $q$, these latter contributions which were not present in Eq. (57) will be suppressed.
@@
where $\mathcal{P}_{Y}\coloneqq\text{id}_{Y}-\mathcal{D}_{Y}$ is a map that projects onto traceless operators on some region $Y$. Defining $\tilde{\Theta}_{i,b_{1}}[\cdot]=\sigma_{i,b_{1}}^{-1/2}\Theta_{i,b_{1}}\sigma_{i,b_{1}}^{-1/2}$ by analogy to Eq. (55), and similarly for the other two maps in (92), our aim is to compute norms $\|\tilde{\Theta}_{i,b_{1}}[X]\|_{2}$ for an arbitrary operator $X$, so we can employ the triangle inequality to bound this by the sum of three contributions corresponding to the three terms in the above expansion.
Using the same steps that led to Eq. (79) (specifically for the case $d^{\prime}=1$, so we can take $L_{0}=\ket{0}\bra{0}$ |), we can compute
$\|\tilde{\Theta}_{i,b_{1}b_{2}}[X]\|_{2}^{2}\leq q^{8}\int{\rm d}\mu_{\rm Br}(U_{b_{1}b_{2}})$
$\times$ $\braket{0}{U_{b_{1}b_{2}}\big{(}(\mathcal{P}_{b_{1}}\otimes\mathcal{P}_{b_{2}})[X]\big{)}U_{b_{1}b_{2}}^{\dagger}}{0}^{2},$ (94)
where the unitary $U_{b_{1}b_{2}}$ acting on the quartet $a=(b_{1},b_{2})$ is no longer sampled from the Haar measure ${\rm d}\mu_{H}$, but rather the distribution obtained from concatenating four Haar-random two-qudit gates in the brickwork arrangement (these correspond to the gates appearing in the third and fourth layers of the original circuit). The analogous expression for $\tilde{\Theta}_{i,b_{1}}$ involves the channel $\mathcal{P}_{b_{1}}$ being replaced by $\mathcal{D}_{b_{1}}$.
With some work, the necessary Haar integrals can be evaluated and we find
@@
where $\tilde{X}=X-\mathcal{D}[X]$ is the traceless part of $X$. Similarly,
@@ Appendix F Proof of advantage over shallow classical circuits
Finally we use Lemma 7 to upper bound the mean error $\overline{\epsilon}$ in the case where $I$ and $J$ are traced out, and $H$ is sent through the distillation channel. Again we can look at a sum over separating domain walls with Boltzmann weights given by $2^{-m|W|/2}$, i.e. the effective inverse temperature is $\beta=\frac{1}{2}m\log 2$. In this case, since $H$, $I$, and $J$ are in the bulk of the system, the relevant domain walls now include loop configurations of the following kinds: 1) a single loop encircling $H$, 2) a single loop encircling $I$ and $J$ together, and 3) two separate loops encircling $I$ and $J$ separately. As we will see, configurations with long loops are strongly suppressed. For a $\sqrt{n}\times\sqrt{n}$ square lattice, we will find it useful to condition on the event that $H$, $I$, and $J$ are all a $\ell_{1}$ distance of at least $\sqrt{n}/4$ from the edges of the system, and also all pairwise separated by at least $\sqrt{n}/4$ (this occurs with constant probability). This effectively allows us to ignore configurations of type 2) above, and also any configurations with domain walls that terminate on the edges of the system (and thus do not form closed loops).